amanda and casey's total bill including tip was $34. Amandas portion of the bill was twice as much as casey's and they each left a $2 tip. how much did each person pay? im stuck on this problem i dont know how to set it up and turn it into an equation i know the answer is amanda payed $20 and casey payed $10 just from logical resoning pkease help to set it up?
Answer by rfer(16322)

you should include their tip
amanda 2x+2
casey x+2
x+2+2x+2=34
3x=30
x=10
x+2=12
2x+2=22